OpenStudy (anonymous):

Find the product of z1 and z2 where z1 = 2(cos 80° + i sin 80°), and z2 = 9(cos 110° + i sin 110°). 11(cos 190° + i sin 190°) 18(cos 8800° + i sin 8800°) 18(cos 190° + i sin 190°) 18(cos 80° + i sin 80°)

jimthompson5910 (jim_thompson5910):

Rule: If \[\Large z_1 = \color{red}{r_1}*\left(\cos(\color{blue}{\theta_1})+i*\sin(\color{blue}{\theta_1})\right)\] \[\Large z_2 = \color{green}{r_2}*\left(\cos(\color{orange}{\theta_2})+i*\sin(\color{orange}{\theta_2})\right)\] ------------------------------------------------------- Then \[\Large z_1*z_2 = \color{red}{r_1}*\color{green}{r_2}*\left(\cos(\color{blue}{\theta_1}+\color{orange}{\theta_2})+i*\sin(\color{blue}{\theta_1}+\color{orange}{\theta_2})\right)\]